# On the Convergence of Decentralized Stochastic Gradient-Tracking with Finite-Time Consensus

### v2

@@ I Introduction
Decentralized strategies to solve problem (1) (see, for example, [sayed2014, cattivelli, ram, extra, next, exact_diff, lian, vlaski_kar_23, pu21, tang18]) rely only on exchanges of information between pairs of neighbors and avoid any centralized processing. They generally consist of two main steps: a local update step, where agents update their model parameters using the gradient or some gradient approximation, and a mixing step, where agents collect and average intermediate quantities (such as models or gradients) with their neighbors via a consensus mechanism:
@@
Performance bounds for iterative decentralized optimization algorithms generally consist of an optimization error and a consensus error. The optimization error arises from the use of iterative gradient-based updates, is common to centralized, federated and decentralized strategies, and represents a lower bound on the performance of distributed optimization algorithms. The consensus error refers to the variation of local models maintained by individual agents, and arises from the limited diffusion of information over the network. It can become negligible for densely or even fully-connected topologies, but for sparsely connected networks, it can contribute significantly to the overall error. Efforts to mitigate the consensus error have included finding the optimal static weights for specific graph topologies [xiao], restricting the optimization to specific graph topologies [song_comm-eff, ying-exp], and using double loop algorithms to enforce consensus [jakovetic, bastianello23].
An alternative approach is the use of finite-time consensus (FTC) sequences [Cesar, kibangou, safavi, sandryhaila]. A sequence of $\tau$ matrices satisfies the finite-time consensus property, if their product equals the scaled all-ones matrix:
@@
FTC sequences arise from the study of linear averaging as a method to achieve consensus amongst agents [ko, Ko_Shi_2009, hendrickx]. Classical weighting rules, such as the Metropolis-Hastings weights [sayed2014], can only guarantee asymptotic convergence, while FTC sequences achieve exact consensus in a finite number of steps. They effectively mimic a fully connected network over $\tau$ steps, where $\tau$ is the consensus number for the graph. The graph diameter serves as a lower bound for $\tau$, while twice the graph’s radius serves as its upper bound [hendrickx].
The advantages of FTC sequences extend beyond the consensus problem and are applicable to more general optimization problems over graphs. FTC sequences have been utilized as the combination weights in decentralized algorithms with demonstrable benefits [Cesar, ying-exp, eusipco]. For instance, FTC sequences have enabled sparser communication between agents without compromising performance [Cesar, ying-exp], while in another work, faster convergence rates were observed for gradient-tracking algorithms using FTC sequences [eusipco].
In practical settings, an exact sequence satisfying (3) may not be available, and instead approximate FTC sequences can be used. Numerical methods for constructing these sequences, such as eigenvalue decompositions [kibangou, safavi, sandryhaila], graph filtering techniques [coutino, segarra], or decentralized learning approaches [eusipco], may introduce inaccuracies. As a result, the generated sequences may only approximate the scaled all-ones matrix. We refer to these as approximate FTC sequences, with the quality of the approximation characterized by $\epsilon_{\tau}$:
@@
Empirical evidence suggests that approximate FTC sequences can nevertheless provide substantial benefit to decentralized optimization [eusipco]. In this work, we analytically quantify the impact of inaccuracies in FTC sequences when used in gradient-tracking schemes.
@@
ATC-Diffusion[yuan20] Strongly Convex Static $\mathcal{O}\left(\frac{\mu\sigma^{2}}{K}+\frac{\mu^{2}\lambda^{2}\sigma^{2}}{1-\lambda}+\frac{\mu^{2}\lambda^{2}\zeta^{2}}{(1-\lambda)^{2}}\right)$ $1-\Theta(\mu\nu)$
ATC-GT[alghunaim] PL-Condition Static $\mathcal{O}\left(\frac{\mu\sigma^{2}}{K}+\frac{\mu^{2}\lambda^{4}\sigma^{2}}{1-\lambda}+\frac{\mu^{4}\lambda^{4}\sigma^{2}}{K(1-\lambda)^{4}}\right)$ $1-\Theta(\mu\nu)$
DIGing[nedic17] Strongly Convex Time-Varying - $\left(1-\Theta(\mu\nu)\right)^{\frac{1}{2\tau}}$
NEXT with FTC[Cesar] Non Convex FTC $\mathcal{O}\left(\frac{\mu\sigma^{2}}{K}+\mu^{2}\tau^{3}\sigma^{2}+\frac{\mu^{4}\tau^{4}\sigma^{2}}{K}\right)$ -
@@
that the noise terms in this work match those of [Cesar].
Any periodic sequence of combination matrices can be considered an approximate FTC sequence and thus our results hold generally for gradient-tracking with time-varying but periodic combination matrices. The same recursion could also be studied in the context of general time-varying matrices (see, for example, [next, nedic16, nedic17, assran19, scutari19, saadatniaki20, huan24]). However, leveraging the periodic structure of the matrices allows us to derive a tighter bound. For example, under the assumption of $\tau$-connectivity (see Assumption 3), linear convergence in [nedic17] is proven with a rate of convergence of $\left(1-\Theta(\mu\nu)\right)^{\frac{1}{2\tau}}$. On the other hand, we prove convergence with an amortized rate of $1-\Theta(\nu\mu)+\Theta(\nu\mu\epsilon_{\tau})$. Our results are summarized in Table LABEL:tab:performance.
We also highlight an interesting phenomenon in the dynamics of the recursion, namely that of a two timescale behavior. The error displays a general downwards trend in periods of $\tau$, however, iterate-on-iterate, the error may grow and shrink, as shown in Fig.s 1 and 4(b). This non-monotonic behavior introduces challenges in the analysis. We therefore derive a bound in terms of the maximum error in periods of $\tau$, in contrast to the usual iterate-on-iterate performance guarantees.
@@ II Analysis
In principle, FTC sequences can be applied in any decentralized algorithm. In this work, we analyze a variant of the gradient-tracking algorithm, Aug-DGM [augdgm]. At each iteration $i$, agent $k$ updates the model variable, $\bm{w}_{k,i}$, and an auxiliary variable that tracks the gradient of the aggregate cost, $\bm{g}_{k,i}$, according to:
@@
As in [Cesar, asilomar], we deviate in our formulation of Aug-DGM by multiplying the gradient approximations by $\mu$ in (5b) rather than $\bm{g}_{\ell,i-1}$ in (5a) as in the original variant of the algorithm [augdgm]. This helps limit the effect of the drift between agents, as discussed in Section III.
@@ II-C Consensus Error
We can compactly describe the dynamics of the consensus error as [Cesar]:
@@ II-E Main Result
Below we combine the consensus error and centroid error bounds from Lemmas (3) and (4) to derive the main result. The key challenge arises from the structure of the recursive dependencies: the centroid error bound at iteration $i$ depends only on the error of the previous iteration, while the consensus error bound depends on the sum of errors across the full FTC sequence from time $m\tau$ to i1i-1. Previous gradient-tracking analyses with perfect FTC ($\epsilon_{\tau}=0$) tackled this by considering the average consensus error terms from iteration zero to $i$ [Cesar]. However, when approximate sequences are employed ($\epsilon_{\tau}>0$), this series will not converge. Instead, we will consider the maximum consensus error within each FTC sequence, resulting in a bound characterizing the maximum error over the $\tau$ iterations of any sequence.
@@ III Simulations and Discussion
Figure 2: MSD for a path graph with 8 agents demonstrating the increasing steady-state error with increasing $\epsilon_{\tau}$. The deterioration in performance is graceful, indicating that Aug-DGM is robust to inaccuracies in the FTC sequence.
@@
Fig. 2 presents the MSD for the optimization problem on a path graph with 16 agents. As before, the MSD is defined as $\mathbb{E}\left\lVert{\bm{\scriptstyle\mathcal{W}}}_{i}-\mathbbm{1}\otimes w^{o}\right\rVert^{2}$}. Three sets of combination matrices are used: one with perfect FTC ($\epsilon_{\tau}=0$) and the remaining two with approximate FTC ($\epsilon_{\tau}=0.3$} and $\epsilon_{\tau}=0.6$} respectively). The weighting rule is provided in the supplementary material. The approximate FTC sequences are generated by adding noise to the non-zero elements in the FTC matrices and adjusting the diagonal entries so that the matrices remain doubly-stochastic. The step size for all three is set to $8\cdot 10^{-3}$}. The results clearly demonstrate the increased steady-state error with higher values of $\epsilon_{\tau}$, corroborating the theoretical prediction that the MSD grows with $\mathcal{O}\left(\frac{\epsilon_{\tau}}{(1-\epsilon_{\tau})^{2}}\right)$}. We note that the restriction of $\epsilon_{\tau}$ to the range [0,0.75]{[0,0.75]} is needed to establish the analytical guarantees, but we have observed in simulations that Aug-DGM can converge even for larger $\epsilon_{\tau}$.
@@
Increased values of $\tau$ also degrades the performance. The MSD in (27a) has a steady-state error that depends on $\mathcal{O}(\mu^{2}\tau^{2})$} which we demonstrate in simulations in Fig. 3. Here, we examine the same optimization problem on different graphs, where each graph has 16 agents with varying values of $\tau$ and with $\mu=5\cdot 10^{-3}$. The results confirm that increasing $\tau$ worsens the performance, consistent with the theoretical predictions.
@@
Finally, we observe an interesting tradeoff between $\tau$ and $\epsilon_{\tau}$. FTC sequences prove especially effective when the consensus number is relatively low, such as in the hypercube in Fig. 4(a) where $K=8,\tau=3$. There is a demonstrable benefit to using the FTC sequence over a classical weighting rule. For graphs with a larger consensus number, the performance degrades. In this case we may choose to set $\tau$ to be smaller than the actual consensus number. The approximate FTC sequence will have a larger approximation error $\epsilon_{\tau}$, but a smaller value of $\tau$. Considering that the MSD depends on the ratio $\frac{\tau^{2}}{(1-\epsilon_{\tau})^{2}}$ this may be beneficial. We demonstrate this in the path graph in Fig. 4(b). The perfect FTC sequence on the path suffers from increasing drift over the course of the sequence. We therefore see the error grow over the sequence, sharply decreasing every $\tau=7$ iterations and demonstrating the two timescale behavior previously described. The best performance is achieved by the approximate FTC case where $\epsilon_{\tau}=0.4$ but $\tau$ is 3, outperforming the perfect FTC case and the metropolis rule, where $\tau=1$ and $\epsilon_{\tau}$ is equal to the second largest eigenvalue, $\lambda=0.95$.
@@ B-B Proof of Lemma (3)
The consensus error is bounded by separating out $\bm{v}_{i}$ on the assumption that the gradient noise is unbiased and pairwise uncorrelated (Assumption (2)) and then using a vector norm bound to separate $\bm{h}_{i}$ as in [koloskova]:
@@
and then using that the spectral norm of a block upper-triangular matrix equals the maximum spectral norm of its diagonal blocks [simovici], in this case $\left\lVert A_{i:m\tau+1}\right\rVert=\epsilon_{\tau}$}.
@@
$\quad+2\tau\mu^{2}\sum_{\mathclap{j=m\tau+1}}^{i}\left(9\beta^{2}\zeta^{2}+9\beta^{2}K\left\lVert\widetilde{\bm{w}}_{c,j-1}\right\rVert^{2}+9\beta^{2}\left\lVert\widehat{{\bm{\scriptstyle\mathcal{W}}}}_{j-1}\right\rVert^{2}+3\sigma^{2}\right)$
@@ Appendix C Centroid Error Relation
We briefly prove Lemma (4), noting that the proof follows arguments in  [chen2015, vlaski21, sayed2014]. We begin by showing that $\frac{1}{K}(\mathbbm{1}^{\mathsf{T}}\otimes I_{M}){\bm{\scriptstyle\mathcal{Y}}}_{i}=0$ from the definition of $y_{i}$ in (31):
@@
where in $(a)$ we added and subtracted $\mu\frac{1}{K}(\mathbbm{1}^{\mathsf{T}}\otimes I_{M})\nabla\mathcal{J}({\bm{\scriptstyle\mathcal{W}}}_{c,i-1})$}, in $(b)$ we used (10e) and in $(c)$ we applied Jensen’s inequality. The first term in (49) is equal to $(1-2\nu\mu+\delta^{2}\mu^{2})\left\lVert\widetilde{\bm{w}}_{c,i-1}\right\rVert^{2}$ (see  [chen2015, vlaski21, sayed2014]). Setting $\alpha=\sqrt{1-2\nu\mu+\delta^{2}\mu^{2}}$:

### v3

@@ I Introduction
Decentralized strategies to solve problem (1) (see, for example, [28, 5, 24, 31, 19, 41, 18, 35, 23, 34]) rely only on exchanges of information between pairs of neighbors and avoid any centralized processing. They generally consist of two main steps: a local update step, where agents update their model parameters using the gradient or some gradient approximation, and a mixing step, where agents collect and average intermediate quantities (such as models or gradients) with their neighbors via a consensus mechanism:
@@
Performance bounds for iterative decentralized optimization algorithms generally consist of an optimization error and a consensus error. The optimization error arises from the use of iterative gradient-based updates, is common to centralized, federated and decentralized strategies, and represents a lower bound on the performance of distributed optimization algorithms. The consensus error refers to the variation of local models maintained by individual agents, and arises from the limited diffusion of information over the network. It can become negligible for densely or even fully-connected topologies, but for sparsely connected networks, it can contribute significantly to the overall error. Efforts to mitigate the consensus error have included finding the optimal static weights for specific graph topologies [37], restricting the optimization to specific graph topologies [33, 39], and using double loop algorithms to enforce consensus [12, 4].
An alternative approach is the use of finite-time consensus (FTC) sequences [22, 13, 26, 27]. A sequence of $\tau$ matrices satisfies the finite-time consensus property, if their product equals the scaled all-ones matrix:
@@
FTC sequences arise from the study of linear averaging as a method to achieve consensus amongst agents [14, 15, 10]. Classical weighting rules, such as the Metropolis-Hastings weights [28], can only guarantee asymptotic convergence, while FTC sequences achieve exact consensus in a finite number of steps. They effectively mimic a fully connected network over $\tau$ steps, where $\tau$ is the consensus number for the graph. The graph diameter serves as a lower bound for $\tau$, while twice the graph’s radius serves as its upper bound [10].
The advantages of FTC sequences extend beyond the consensus problem and are applicable to more general optimization problems over graphs. FTC sequences have been utilized as the combination weights in decentralized algorithms with demonstrable benefits [22, 39, 8]. For instance, FTC sequences have enabled sparser communication between agents without compromising performance [22, 39], while in another work, faster convergence rates were observed for gradient-tracking algorithms using FTC sequences [8].
In practical settings, an exact sequence satisfying (3) may not be available, and instead approximate FTC sequences can be used. Numerical methods for constructing these sequences, such as eigenvalue decompositions [13, 26, 27], graph filtering techniques [7, 30], or decentralized learning approaches [8], may introduce inaccuracies. As a result, the generated sequences may only approximate the scaled all-ones matrix. We refer to these as approximate FTC sequences, with the quality of the approximation characterized by $\epsilon_{\tau}$:
@@
Empirical evidence suggests that approximate FTC sequences can nevertheless provide substantial benefit to decentralized optimization [8]. We demonstrate this in Fig. 1 for a decentralized linear least squares problem, with simulation details provided in Sec. III. For the considered graph, no analytical FTC sequence is known. An approximate sequence can be obtained using the numerical method in [8] ($\epsilon_{\tau}=0.7$$\tau=4$), which we compare with the fastest distributed linear averaging (FDLA) matrix [37], a static combination matrix designed to have the largest mixing rate between agents. Gradient-tracking performance with the approximate sequence clearly outperforms the FDLA case, demonstrating that approximate FTC sequences can improve performance despite not achieving exact consensus.
Since exact FTC sequences are only known for a handful of graph families, and approximate sequences demonstrate empirical benefits, these observations motivate a deeper analysis into their behavior. In this work, we analytically quantify the impact of inaccuracies in FTC sequences when used in gradient-tracking schemes, highlighting both the benefits and trade-offs of their use.
@@
ATC-Diffusion[40] Strongly Convex Static $\mathcal{O}\left(\frac{\mu\sigma^{2}}{K}+\frac{\mu^{2}\lambda^{2}\sigma^{2}}{1-\lambda}+\frac{\mu^{2}\lambda^{2}\zeta^{2}}{(1-\lambda)^{2}}\right)$ $1-\Theta(\mu\nu)$
ATC-GT[2] PL-Condition Static $\mathcal{O}\left(\frac{\mu\sigma^{2}}{K}+\frac{\mu^{2}\lambda^{4}\sigma^{2}}{1-\lambda}+\frac{\mu^{4}\lambda^{4}\sigma^{2}}{K(1-\lambda)^{4}}\right)$ $1-\Theta(\mu\nu)$
DIGing[20] Strongly Convex Time-Varying - $\left(1-\Theta(\mu\nu)\right)^{\frac{1}{2\tau}}$
NEXT with FTC[22] Non Convex FTC $\mathcal{O}\left(\frac{\mu\sigma^{2}}{K}+\mu^{2}\tau^{3}\sigma^{2}+\frac{\mu^{4}\tau^{4}\sigma^{2}}{K}\right)$ -
@@
that the noise terms in this work match those of [22].
Any periodic sequence of combination matrices can be considered an approximate FTC sequence and thus our results hold generally for gradient-tracking with time-varying but periodic combination matrices. The same recursion could also be studied in the context of general time-varying matrices (see, for example, [19, 21, 20, 3, 29, 25, 17]). However, leveraging the periodic structure of the matrices allows us to derive a tighter bound. For example, under the assumption of $\tau$-connectivity (see Assumption 3), linear convergence in [20] is proven with a rate of convergence of $\left(1-\Theta(\mu\nu)\right)^{\frac{1}{2\tau}}$. On the other hand, we prove convergence with an amortized rate of $1-\Theta(\nu\mu)+\Theta(\nu\mu\epsilon_{\tau})$, which is faster by a factor of $\tau$. Our results are summarized in Table LABEL:tab:performance.
We also highlight an interesting phenomenon in the dynamics of the recursion, namely that of a two timescale behavior. The error displays a general downwards trend in periods of $\tau$, however, iterate-on-iterate, the error may grow and shrink, as shown in Fig.s 2 and 5(b). This non-monotonic behavior introduces challenges in the analysis. We therefore derive a bound in terms of the maximum error in periods of $\tau$, in contrast to the usual iterate-on-iterate performance guarantees.
@@
(a) Graph
(b) Performance
Figure 1: The performance of gradient-tracking can be improved using approximate FTC sequences. Here we compare with the fastest distributed linear averaging (FDLA) matrix [37].
@@ II Analysis
In principle, FTC sequences can be applied in any decentralized algorithm. In this work, we analyze a variant of the gradient-tracking algorithm, Aug-DGM [38]. At each iteration $i$, agent $k$ updates the model variable, $\bm{w}_{k,i}$, and an auxiliary variable that tracks the gradient of the aggregate cost, $\bm{g}_{k,i}$, according to:
@@
As in [22, 9], we deviate in our formulation of Aug-DGM by multiplying the gradient approximations by $\mu$ in (5b) rather than $\bm{g}_{\ell,i-1}$ in (5a) as in the original variant of the algorithm [38]. This helps limit the effect of the drift between agents, as discussed in Section III.
@@ II-C Consensus Error
We can compactly describe the dynamics of the consensus error as [22]:
@@ II-E Main Result
Below we combine the consensus error and centroid error bounds from Lemmas (3) and (4) to derive the main result. The key challenge arises from the structure of the recursive dependencies: the centroid error bound at iteration $i$ depends only on the error of the previous iteration, while the consensus error bound depends on the sum of errors across the full FTC sequence from time $m\tau$ to i1i-1. Previous gradient-tracking analyses with perfect FTC ($\epsilon_{\tau}=0$) tackled this by considering the average consensus error terms from iteration zero to $i$ [22]. However, when approximate sequences are employed ($\epsilon_{\tau}>0$), this series will not converge. Instead, we will consider the maximum consensus error within each FTC sequence, resulting in a bound characterizing the maximum error over the $\tau$ iterations of any sequence.
@@ III Simulations and Discussion
Figure 3: Steady-state error against $\epsilon_{\tau}$ for randomly generated graphs (with $\tau$ fixed to 5) demonstrating the deterioration in performance with increasing values of $\epsilon_{\tau}$. To simplify comparison of the trend the theoretical prediction has been shifted downwards.
@@
Fig. 3 demonstrates the steady-state MSD against $\epsilon_{\tau}$ for the optimization problem on randomly generated 8-node graphs. The graphs are sampled from three families: Erdős–Rényi (binomial) graphs with an edge probability $p\sim U(0.1,0.9)$}, Barabasi-Albert graphs [1] with edge numbers $m\sim U(1,6)$}, or random 8-node trees. Graph connectivity is verified, otherwise a new graph is sampled. The approximate FTC matrices are then generated using the iterative method described in [8] with $\tau$ is set to 5. The resulting values of $\epsilon_{\tau}$, vary depending on the solver’s progress and whether the graph’s actual consensus number is 5. This setup provides a realistic example of the numerical errors typically associated with generating FTC matrices. These learned approximate FTC matrices are then used in the Aug-DGM algorithm with stochastic gradients and $\mu=0.01$. In total, 1000 independent simulations were performed. The results in Fig. 3 clearly demonstrate the increased steady-state error with higher values of $\epsilon_{\tau}$, corroborating the theoretical prediction that the MSD is proportional to $\mathcal{O}\left(\frac{1}{(1-\epsilon_{\tau})^{2}}\right)$}. We note that the restriction of $\epsilon_{\tau}$ to the range [0,0.75]{[0,0.75]} is needed to establish the analytical guarantees, but we observe in simulations that Aug-DGM can converge even for larger $\epsilon_{\tau}$.
@@
Increased values of $\tau$ also degrade the performance. The MSD in (27a) has a steady-state error that depends on $\mathcal{O}(\mu^{2}\tau^{2})$} which we demonstrate in Fig. 4. Here, we examine the same least-squares optimization problem on different 16-node graphs, where the exact FTC sequences are known ($\epsilon_{\tau}=0)$: namely, a 16-hypercube [11], $4\times 4$ square-grid graph, 16-cycle, and 16-path. The FTC sequences for the cycle and path graphs are provided in the supplementary material. The grid graph’s sequence is the concatenation of the FTC sequence for the 4-node paths acting first along the horizontal and then the vertical dimensions (leveraging the fact that the $4\times 4$ square-grid graph is the Cartesian product of two 4-paths). The results in (27a) confirm that increasing $\tau$ worsens the performance, consistent with the theoretical predictions.
@@
Finally, we observe an interesting trade-off between $\tau$ and $\epsilon_{\tau}$. Exact FTC sequences prove especially effective when the consensus number is relatively low, such as in the hypercube in Fig. 5(a) where $K=8,\tau=3$. There is a demonstrable benefit to using the exact FTC sequence over a classical weighting rule. For graphs with a larger consensus number, the performance degrades. In this case we may choose to set $\tau$ to be smaller than the actual consensus number. The approximate FTC sequence will have a larger approximation error $\epsilon_{\tau}$, but a smaller value of $\tau$. Considering that the MSD depends on the ratio $\frac{\tau^{2}}{(1-\epsilon_{\tau})^{2}}$ this may be beneficial. We demonstrate this in the path graph in Fig. 5(b). The perfect FTC sequence on the path suffers from increasing drift over the course of the sequence. We therefore see the error grow over the sequence, sharply decreasing every $\tau=7$ iterations and demonstrating the two timescale behavior previously described. The best performance is achieved by the approximate FTC case ($\epsilon_{\tau}=0.4$). Matrices are learned using the method in [8] with $\tau=3$. The approximate case outperforms both the perfect FTC case and the static metropolis rule. More generally we can conclude from our theoretical predictions and empirical results that approximate FTC sequences provide a benefit over a static weighting rule (with second largest eigenvalue $\lambda$) when $\frac{\tau^{2}}{(1-\epsilon_{\tau})^{2}}$ is small relative to $\frac{1}{(1-\lambda)^{2}}$.
@@ B-B Proof of Lemma (3)
The consensus error is bounded by separating out $\bm{v}_{i}$ on the assumption that the gradient noise is unbiased and pairwise uncorrelated (Assumption (2)) and then using a vector norm bound to separate $\bm{h}_{i}$ as in [16]:
@@
and then using that the spectral norm of a block upper-triangular matrix equals the maximum spectral norm of its diagonal blocks [32], in this case $\left\lVert A_{i:m\tau+1}\right\rVert=\epsilon_{\tau}$}.
@@
$\quad+2\tau\mu^{2}\sum_{\mathclap{j=m\tau+1}}^{i}\big(9\beta^{2}\zeta^{2}+9\beta^{2}K\left\lVert\widetilde{\bm{w}}_{c,j-1}\right\rVert^{2}$
$\qquad+9\beta^{2}\left\lVert\widehat{{\bm{\scriptstyle\mathcal{W}}}}_{j-1}\right\rVert^{2}+3\sigma^{2}\big)$
@@ Appendix C Centroid Error Relation
We briefly prove Lemma (4), noting that the proof follows arguments in  [6, 36, 28]. We begin by showing that $\frac{1}{K}(\mathbbm{1}^{\mathsf{T}}\otimes I_{M}){\bm{\scriptstyle\mathcal{Y}}}_{i}=0$ from the definition of $y_{i}$ in (31):
@@
where in $(a)$ we added and subtracted $\mu\frac{1}{K}(\mathbbm{1}^{\mathsf{T}}\otimes I_{M})\nabla\mathcal{J}({\bm{\scriptstyle\mathcal{W}}}_{c,i-1})$}, in $(b)$ we used (10e) and in $(c)$ we applied Jensen’s inequality. The first term in (49) is equal to $(1-2\nu\mu+\delta^{2}\mu^{2})\left\lVert\widetilde{\bm{w}}_{c,i-1}\right\rVert^{2}$ (see  [6, 36, 28]). Setting $\alpha=\sqrt{1-2\nu\mu+\delta^{2}\mu^{2}}$: